OpenStudy (anonymous):

2log-a^x + log-a^2 =log-a (5x + 3) solve for x?

OpenStudy (anonymous):

it's # 5

OpenStudy (anonymous):

Sorry idk :(

OpenStudy (anonymous):

Use a calculator the Log function. I forgot how to do this

OpenStudy (anonymous):

aw it's ok.

jhonyy9 (jhonyy9):

so log_a what mean ? logarithm base how many ?

OpenStudy (anonymous):

i'm not sure

jhonyy9 (jhonyy9):

so on the attached page i see it clearly there is log base a

jhonyy9 (jhonyy9):

so what you have wrote above here is the 59 exercise yes ?

OpenStudy (anonymous):

I didn't write any of it, my teacher did. :/

jhonyy9 (jhonyy9):

ok do you know the property of logarithms ?

jhonyy9 (jhonyy9):

for example log a + log b = ?

OpenStudy (anonymous):

I'm not so sure? you mean like when you add them log a + log B = log AB and that stuff?

jhonyy9 (jhonyy9):

yes you are right

jhonyy9 (jhonyy9):

and do you know how will be log x^2 = ?

jhonyy9 (jhonyy9):

this log mean always base 2

OpenStudy (anonymous):

which one are you looking at? I'm looking at #6 and I think the base is 5.

OpenStudy (anonymous):

oh whoops wait

OpenStudy (anonymous):

no you're right I was looking at the wrong one

OpenStudy (anonymous):

ok so the base is 2, awesome

jhonyy9 (jhonyy9):

check the 5 nr.

OpenStudy (anonymous):

with that knowledge, should I go ahead and add them together?

OpenStudy (anonymous):

but how do I know the base is 2? ahhh I'm so sorry I'm so confused.

jhonyy9 (jhonyy9):

ok just a moment from these 3 what you need to solve it ?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

and in the last one, t.

OpenStudy (anonymous):

she gave us the answers so that we can know if we did it correctly but I still have to do the work.

jhonyy9 (jhonyy9):

can you rewriting it here now ?

jhonyy9 (jhonyy9):

2loga x = ?

jhonyy9 (jhonyy9):

this 2 what mean ?

OpenStudy (anonymous):

ohh it's the exponent.

OpenStudy (anonymous):

a squared

jhonyy9 (jhonyy9):

not sure is the exponent of x because a is the base of logarithm

OpenStudy (anonymous):

Well that's what I think... That's just how she wrote it.

jhonyy9 (jhonyy9):

ok so than can you rewriting this exercise what you like we solve it together and and the second will solve it you ok ?

OpenStudy (anonymous):

ok sounds good.

jhonyy9 (jhonyy9):

what you wann we solving here together ?

OpenStudy (anonymous):

so 2loga^x + loga^2 = loga(5x+3)

OpenStudy (anonymous):

sure. Like that?

jhonyy9 (jhonyy9):

ATTENTION never not will be right using the sign of exponent how you have wrote above because this mean ,different thing here what you have wrote there mean 2 time logarithm base 2 of a on exponent x

jhonyy9 (jhonyy9):

you can writing logarithm base a of x like log_a x

jhonyy9 (jhonyy9):

rewrit it please so

OpenStudy (anonymous):

ok so 2loga_x + loga_2 = loga (5x+3)

jhonyy9 (jhonyy9):

the _ sign you need writing next a so log_a x ok ?

jhonyy9 (jhonyy9):

hence 2log_a x +log_a 2 = log_a (5x+3) so than now on the left part what you can doing ?

OpenStudy (anonymous):

you can combine the two logs.

jhonyy9 (jhonyy9):

yes how ?

OpenStudy (anonymous):

you add them. so would it be log_a x2 = log_a(5x+3)? I'm just guessing.

jhonyy9 (jhonyy9):

ok but where is the 2 from the 2log_a x ?

OpenStudy (anonymous):

I'm not sure. All I can think to do is put it with the x.

jhonyy9 (jhonyy9):

you have wrote it above that the 2 from 2log_a x what mean ?

jhonyy9 (jhonyy9):

check it please above on your first answers

OpenStudy (anonymous):

I'm sorry I have to go, I think I'll just ask my teacher tomorrow. But thank you for your time!

jhonyy9 (jhonyy9):

2log_a x = log_a x^2 yes ?

jhonyy9 (jhonyy9):

than there will be log_a x^2 + log_a 2 = log_a 2x^2 ok ?